4th term Of an arithmetic sequence is 17. Sum of first 20 terms is 990. Find first term and common difference.
Answer by mananth(16946)

4th term Of an arithmetic sequence is 17. Sum of first 20 terms is 990. Find first term and common difference.
4th term Of an arithmetic sequence is 17.
t4 =17
a+3d=17.................(1)
Sum of first 20 terms is 990.
S20 = 990
Sn = n/2(2a+(n-1)d)
20/2(2a+(20-1)d)=990
10(2a+19d)=990
2a+19d =99.................(2)
Sole (1) & (2)
a+3d=17
multiply by -2
-2a-6d=-34
add this equation to (2)
13d=65
d=5
a+3d=17
a+15=17
a=2
2,7,12,17,.........


Find first term and common difference.